OpenStudy (anonymous):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for three different years: Number of years 1 2 3 Option 1 (amount in dollars) 1100 1200 1300 Option 2 (amount in dollars) 1100 1210 1331 Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option.

OpenStudy (anonymous):

F(n) = 100 ^ n + 1000 F(n) = 1000 * 1.1^n

OpenStudy (dannyo19):

So there's a 1 dollar difference. That could buy you a delicious avocado. I'd say that's significant, but seems to be a matter of opinion.

OpenStudy (anonymous):

F(20) = 100 ^ n + 1000 F(20) = 1000 * 1.1^n

OpenStudy (elvisiscool69):

belinda is balling with money